The tens digit of a two-digit number is 1 more the twice the units digit. The number is 26 times the units digit. Find the number.
Answer by Nate(3500)

Number: 10a + b
a = 1 + 2b
and
10a + b = 26b
10(1 + 2b) = 25b
10 + 20b = 25b
10 = 5b
2 = b
Plug:
a = 1 + 2b = 1 + 4 = 5
Number: 52